# A Damage-Driven Model for Duchenne Muscular Dystrophy: Early-Stage Dynamics and Invasion Thresholds

### 31 Mar 2026 v2

@@ 4.3 Uniform-in-time estimates
$\upsilon_{\tau}\nabla m=\int_{\tau}^{t}e^{-D_{m}\Delta(t-\tau-s)}\nabla\left(\upsilon_{\tau}\nu(d-m)-\upsilon_{\tau}{{}^{\prime}}m-\nabla\cdot\left(\chi(m,c)\nabla c\right)\right)ds\,.$ (4.83)
Using the $L^{p}-L^{q}$ estimate (see [13]) of the heat kernel, we have
$\sup_{\tau<t<\tau+2}\|\upsilon_{\tau}m\|_{L^{\infty}(\Omega)}\leq\overline{C}^{\ast}\left(\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}+\|d\|_{L^{p}(\Omega_{\tau,\tau+2})}\right)\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-n/2p}\right)ds+\\ +\overline{C}^{\ast}\chi_{0}\|\nabla\,c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-\frac{1}{2}-\frac{n}{2p}}\right)ds\,$, (4.84)
@@
$\sup_{\tau<t<\tau+2}\|\upsilon_{\tau}\nabla m\|_{L^{\infty}(\Omega)}\leq\overline{C}^{\ast}\left(\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}+\|d\|_{L^{p}(\Omega_{\tau,\tau+2})}\right)\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-\frac{1}{2}-\frac{n}{2p}}\right)ds+\\ +\overline{C}^{\ast}\chi_{0}\|\nabla\,c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-1-\frac{n}{2p}}\right)ds\,$, (4.85)
where $\overline{C}^{\ast}$ depends on $\Omega$, $n$, and the parameters of the system, but not on $\tau$.
Choosing $p>n$, and using (4.70) and (4.71) together with the properties of $\upsilon_{\tau}$, we obtain (4.77) and (4.78) for $m$.
Proposition 15, Proposition 16 and Proposition 20 complete the proof of Theorem 5.
@@ 7 Discussion and Conclusions
The present work also highlights several limitations and open problems. The existence and stability properties of the diseased equilibrium remain difficult to characterize in full generality, due to the nonlinear coupling and the large number of parameters. Nevertheless, in a parameter regime consistent with the early-stage scenario considered here, the pathological equilibrium can be partially characterized and its existence is directly linked to the invasion threshold. A more detailed bifurcation analysis with respect to biologically relevant parameters could provide deeper insight into long-term disease outcomes. Moreover, the observation that invasion may occur even in the absence of a stable pathological equilibrium suggests that disease progression can be driven by dynamical mechanisms that do not rely on convergence to a steady state. Extending the analysis beyond the early-stage regime, particularly in the vicinity of pathological equilibria, is a natural next step. Numerical evidence indicates the presence of bistable regimes in which healthy and pathological states coexist, and a deeper examination of the associated stability and bifurcation structure could further clarify the mechanisms governing long-term disease progression.

### 03 Apr 2026 v3

@@ 4.3 Uniform-in-time estimates
$\upsilon_{\tau}\nabla m=\int_{\tau}^{t}e^{-D_{m}\Delta(t-\tau-s)}\nabla\left(\upsilon_{\tau}\nu(d-m)+\upsilon_{\tau}{{}^{\prime}}m-\upsilon_{\tau}\nabla\cdot\left(\chi(m,c)\nabla c\right)\right)ds\,.$ (4.77)
Using the $L^{p}-L^{q}$ estimate (see [14]) of the heat kernel, we have:
$\sup_{\tau<t<\tau+2}$ $\|\upsilon_{\tau}m\|_{L^{\infty}(\Omega)}\leq$
$\overline{C}^{\ast}\left(\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}+\|d\|_{L^{p}(\Omega_{\tau,\tau+2})}\right)\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-n/2p}\right)ds+$
$+\overline{C}^{\ast}\chi_{0}\|\nabla\,c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-\frac{1}{2}-\frac{n}{2p}}\right)ds\,$, (4.78)
@@
$\sup_{\tau<t<\tau+2}$ $\|\upsilon_{\tau}\nabla m\|_{L^{\infty}(\Omega)}\leq$
$\overline{C}^{\ast}\left(\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}+\|d\|_{L^{p}(\Omega_{\tau,\tau+2})}\right)\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-\frac{1}{2}-\frac{n}{2p}}\right)ds+$
$+\overline{C}^{\ast}\chi_{0}\|\nabla\,c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\|m\|_{L^{p}(\Omega_{\tau,\tau+2})}\int_{\tau}^{\tau+2}\left(1+(t-\tau-s)^{-1-\frac{n}{2p}}\right)ds\,$, (4.79)
where $\overline{C}^{\ast}$ depends on $\Omega$, $n$ and the parameters of the system, but not on $\tau$.
Choosing $p>n$ and using (4.66), (4.67) together with the properties of $\upsilon_{\tau}$, we obtain (4.73) and (4.74) for $m$.   Proposition 15, Proposition 16 and Proposition 20 complete the proof of Theorem 5.
@@ 7 Discussion and Conclusions
The present work also highlights several limitations and open problems. The existence and stability properties of the diseased equilibrium remain difficult to characterize in full generality, due to the nonlinear coupling and the large number of parameters. Nevertheless, in a parameter regime consistent with the early-stage scenario considered here, the pathological equilibrium can be partially characterized and its existence is directly linked to the invasion threshold. A more detailed bifurcation analysis with respect to biologically relevant parameters could provide deeper insight into long-term disease outcomes. Extending the analysis beyond the early-stage regime, particularly in the vicinity of pathological equilibria, is a natural next step. Numerical evidence indicates the presence of bistable regimes in which healthy and pathological states coexist, and a deeper examination of the associated stability and bifurcation structure could further clarify the mechanisms governing long-term disease progression.